solve by the substitution method.
8x+7y=9
-3x+y=22

Answer by Gogonati(855)

From the second equation isolate y: y=3x+22. Substitute this value for y in the first equation:8x+7(3x+22)=9, solving this linear equation we get:29x=-145, and
final x=-145/29. Plugging x=-145/29 in the equation y=3x+22 we get: y=3(-145/29)+22,
and final y=7.
Answer: the solution of our system is the ordered pair (-145/29, 7).